OpenStudy (anonymous):

show that the square of any odd integer is of the form 4q+1 for some integer q.

OpenStudy (shubhamsrg):

any odd integer is in the form of (2n +1),,where n is some integer sq of it = 4n^2 + 1 + 4n = 4(n^2 +n) + 1 = 4q + 1 where q=(n^2 +n) hence proved.. hmmn

OpenStudy (kinggeorge):

On a related tangent, the square of any odd integer is also of the form 8k+1 for some integer k. This is a slightly more involved proof (but only slightly).

OpenStudy (anonymous):

can any1 give step by step?

OpenStudy (kinggeorge):

@shubhamsrg already gave the step by step for the 4q+1 problem. If you want one for the 8k+1 problem, I'll be happy to elaborate.

OpenStudy (shubhamsrg):

induction may also be used here as a successful tool..

OpenStudy (shubhamsrg):

both for 4k+1 and 8k+1

OpenStudy (shubhamsrg):

and for that 8k+1 by my method,,we know n^2 + n = n(n+1) which is always even (ofcorse) = 2k,for some integer k.. hence it takes the shape of 8k+1 is there any other approach than this @KingGeorge sir ?

OpenStudy (kinggeorge):

There are a few ways to approach the 8k+1 problem. Simplest, is modular arithmetic. Next simplest, is the way you just described. More complicated, but still straightforward, is to look at \(4n^2+4n+1\) and look at the case where \(n\) is even vs. \(n\) is odd. There's also induction, as you also mentioned.

OpenStudy (shubhamsrg):

yeah i see what you're intending to explain :)

mathslover (mathslover):

@jyosnika shubhamsrg has the right answer ...